whats the equation for f(x)=x^2 is vertically stretched by a factor of 9 and translated 2 units to create g
Answer by stanbon(75887)

whats the equation for f(x)=x^2 is vertically stretched by a factor of 9 and translated 2 units to create g
-----
If it is translated to the right you
would have g(x) = 9(x-2)^2
------
If it is translated to the left you
would have g(x) = 9(x+2)^2
===============================
Cheers,
Stan H.